# When the Rules Change: Adaptive Signal Extraction via Kalman Filtering and Markov-Switching Regimes

### v1

@@ 7 Discussion and Conclusion
7.1 Summary of Findings
This paper develops a dynamic, state-dependent framework for analyzing order flow–return relationships in the Korean stock market. Our three-pillar approach—Adaptive Kalman Filtering, Markov-Switching Regimes, and Asymmetric Response Functions—reveals systematic patterns that static models obscure.
The central finding is that foreign investor predictive power increases dramatically during crisis periods. The regime-conditional coefficient rises from 0.00023 in Bull markets to 0.00204 in Crisis periods, an 8.9-fold increase. This suggests that either (a) foreign investors possess superior information that becomes more valuable during stress, or (b) noise trading by retail investors increases during crises, making the relative signal strength of foreign flow more pronounced.
Individual investors exhibit momentum-chasing behavior, with response coefficients to positive shocks exceeding those to negative shocks by a factor of 6.3. This asymmetry is consistent with behavioral biases documented in the literature and contributes to understanding why retail participation often amplifies rather than dampens market volatility.
@@ 7.2 Implications
For Practitioners: The methodology provides a framework for adaptive signal processing that automatically discounts noisy observations during stress periods. While overall sample performance is challenging, the strong results during the COVID crisis period (Sharpe 1.08) demonstrate the value of regime-conditional positioning during extreme events.
@@ 7.4 Conclusion
Static models of market microstructure assume stable relationships that empirically fail during stress periods—precisely when accurate models are most needed. This paper demonstrates that explicitly modeling regime-dependence and asymmetric responses yields richer understanding of how different investor classes interact with market conditions.
The paper’s primary contribution is methodological: demonstrating how state-space techniques can be adapted to capture dynamic relationships in financial markets. The empirical findings—particularly the 8.9-fold increase in foreign investor predictive power during crises—provide new evidence on information heterogeneity and its market-condition dependence.
Future research might extend this framework to higher-frequency data, investigate the sources of foreign investor information advantage, or apply similar techniques to other markets with investor-type classification. Understanding when and why order flow relationships change remains central to both academic research and practical market analysis.

### v2

@@ 7 Discussion and Conclusion
7.1 Synthesis
Taken together, the three methodological pillars converge on a single structural conclusion: information asymmetry in the Korean market is not a constant—it is regime-dependent, amplified by the simultaneous increase in both informed activity and uninformed noise during stress. The Kalman filter reveals that raw signals become noisier in high-volatility periods, yet the regime-switching model shows that the signal-to-noise ratio for foreign flow actually improves during crises. The asymmetric response analysis resolves this apparent paradox: crisis episodes trigger opposing behavioral responses—foreign contrarian buying and retail momentum-chasing—that widen the informational wedge between investor types precisely when volatility is highest.
The information-theoretic validation corroborates this dual-driver mechanism independently. Viterbi-decoded “Informed” trading days concentrate disproportionately in crisis regimes, and foreign flow is predictive only on those days, confirming that elevated informed activity and elevated noise co-occur during stress. This distinguishes our finding from simpler accounts in which one side alone—either superior foreign information or inferior retail behavior—explains the regime dependence. However, the decisive out-of-sample failure (Section 5.5) demonstrates that even a well-characterized in-sample regularity need not generalize, likely because the specific mix of informed and noise intensity shifts across macroeconomic episodes.
@@ 7.2 Implications
For Practitioners: The methodology provides a framework for adaptive signal processing that automatically discounts noisy observations during stress periods. While overall sample performance is challenging, the strong crisis-period results (Section 5.4) demonstrate the value of regime-conditional positioning during extreme events.
@@ 7.4 Conclusion
The central lesson of this paper is that the informational role of different investor types is not a fixed quantity—it shifts with market conditions, concentrating in crisis periods and attenuating during calm ones. Static coefficients obscure this variation, and models calibrated in one regime provide poor guidance in another.
Our evidence for this claim rests on three mutually reinforcing pillars: parametric regime-switching analysis that quantifies the shift (Section 5.2), behavioral asymmetry estimation that documents the opposing investor-type reactions driving it (Section 5.3), and non-parametric information-theoretic validation that independently corroborates the regime structure (Section 6.2). The convergence across methods strengthens confidence in the descriptive finding while the out-of-sample failure (Section 5.5) honestly delimits its predictive reach.
Future research might investigate whether similar regime-dependent patterns hold in other emerging markets with investor-type classification, explore the microeconomic sources of foreign investor information advantage during crises, or extend the analysis to higher-frequency data. Understanding how and why order flow relationships vary across market conditions remains central to both academic microstructure research and practical market analysis.
@@ Appendix A Additional Figures
Figure 12: Rolling Jensen-Shannon Divergence. Top panel: JSD of $S^{TV}$ and $S^{MC}$ distributions against their respective reference distributions. Bottom panel: cross-JSD between the $S^{TV}$ and $S^{MC}$ distributions over time. Dashed red line indicates the 0.1 nats regime-shift threshold.
Figure 13: JSD-Based Adaptive Blending Weights. Top panel: time-varying weights $w^{TV}_{t}$ and $w^{MC}_{t}$ (faint red vertical lines indicate Crisis regime days). Bottom panel: resulting $S^{Adaptive}$ signal versus static $S^{TV}$.
Figure 14: Viterbi HMM State Assignment. Top: foreign signal colored by decoded state (green = Informed, red = Noise). Middle: binary state assignment over time. Bottom: market return for reference.
Figure 15: Combined IT Validation Summary. Left: predictive $t$-statistics by signal and day-type subset. Right: regression coefficients with 95% block-bootstrap confidence intervals.